2) A class has 25 students. Each day, three students are chosen to present their work at the board. In total, how many such 3 student groups are possible?

3) President Trump’s approval rating is 42%. Suppose that 10 people were chosen at random a) Find the probability that 5 of the 10 people approve of the job President Trump is doing.

b) Find the probability that at most 3 of the 10 people approve of the job President Trump is doing.

c) Find the probability that at least 3 of 10 people approve of the job President Trump is doing.
Answer #1

3)

Sample size , n =    10
Probability of an event of interest, p =   0.42

a)

P ( X =    5   ) = C(10,5) * 0.42^5 * (1-0.42)^5 =            0.2162 (answer)

b)

P(at most 3) = P(X≤3) = P(X=0) + P(X=1) + P(X=2) + P(X=3) =0.0043+0.0312+0.1017+0.1963 = 0.3335

c)

P(at least 3) = 1 - P(X<3) = 1 - [ P(X=0) + P(X=1) + P(X=2) ] = 0.8628